# LPR: Large Language Models-Aided Program Reduction

### v2

@@ Abstract.
Our thorough evaluation on 50 benchmarks across three programming languages (i.e., C, Rust and JavaScript) has demonstrated LPR’s practicality and superiority over Vulcan, the state-of-the-art language-generic program reducer. For effectiveness, LPR surpasses Vulcan by producing 24.93%, 4.47%, and 11.71% smaller programs on benchmarks in C, Rust and JavaScript, separately. Moreover, LPR and Vulcan have the potential to complement each other. For the C language for which C-Reduce is optimized, by applying Vulcan to the output produced by LPR, we can attain program sizes that are on par with those achieved by C-Reduce. For efficiency, LPR is more efficient when reducing large and complex programs, taking %, %, % less time than Vulcan to finish all the benchmarks in C, Rust and JavaScript, separately.
@@ 1. Introduction
Program reduction techniques (Sun et al., 2018; Xu et al., 2023; Regehr et al., 2012; Kalhauge and Palsberg, 2019, 2021; Niemetz and Biere, 2013; Tian et al., 2023b; Wang et al., 2021; Misherghi and Su, 2006; Donaldson et al., 2021) aim to facilitate compiler debugging by minimizing the bug-triggering programs with efficacy and efficiency. Given a program and a property that preserves, program reduction techniques (a.k.a., program reducers) produce a minimal program that still preserves . Program reduction has been widely used in various software engineering tasks (Donaldson and MacIver, 2021), especially in compiler testing and debugging (Livinskii et al., 2020; Lecoeur et al., 2023).
However, a critical challenge in program reduction has not been properly addressed, i.e., the trade-off between generality across languages and specificity to a certain language. Currently, there are two categories of program reduction techniques: language-specific (Regehr et al., 2012; Kalhauge and Palsberg, 2019, 2021) and language-generic (Zeller and Hildebrandt, 2002; Misherghi and Su, 2006; Sun et al., 2018; Xu et al., 2023; Zhang et al., 2023). The former category leverages language-specific semantics to transform and shrink programs in certain languages, while the latter only uses transformations applicable to any programming language. Although language-specific reducers are usually more effective in reduction, designing an effective reducer for a specific language, especially designing language-specific transformations, requires a deep understanding of language features and a significant amount of time and engineering effort. Therefore, only a limited set of languages have custom-designed reducers, such as C (Regehr et al., 2012), Java (Kalhauge and Palsberg, 2019, 2021), and SMT-LIBv2 (Niemetz and Biere, 2013). Meanwhile, language-generic reducers can be applied to diverse languages, but lack the knowledge of language features and semantics and thus are incapable of performing language-specific transformations (e.g., function inlining) that can enable further reduction. As a result, they cannot utilize peculiar features of each language to achieve optimal reduced programs.
@@
Meanwhile, we also notice that recent progress in Large Language Models (LLMs) could be a powerful assistant in performing language-specific transformations, like its performance in other scenarios of software engineering tasks such as code generation and test generation (Sun et al., 2023; Jain et al., 2022; Wei et al., 2023; Deng et al., 2023a; Fan et al., 2023; Xia et al., 2023b; Gu, 2023). Specifically, LLMs are trained with massive programs, and they have started to exhibit the ability to analyze and transform programs in prevalent languages. If we can properly leverage this ability for program reduction, we may have a language-generic reducer being aware of the semantics of various prevalent languages. Besides, the utilization of LLMs can simplify the customization and extension of reducers, as it would be time-consuming and labor-intensive to manually implement a language-specific reducer or add functionality to an existing one (such as C-Reduce which uses the Clang frontend to implement C-specific program transformations).
Challenges of Using LLMs. LLMs are not the silver bullet to program reduction. They face challenges like understanding subtle differences (Li et al., 2023) in code, getting easily distracted by irrelevant context (Shi et al., 2023) and performing poorly with large inputs due to catastrophic forgetting (Kemker et al., 2018; Arora et al., 2019; Goodfellow et al., 2013). Specifically, in program reduction, LLMs may not be capable of reducing large programs directly because of distraction and catastrophic forgetting. Without effective guidance, LLMs are unclear about what transformations to perform.
LLMs-Aided Program Reduction (LPR). We propose LPR (Large language model aided program reduction) in this paper, which is, to the best of our knowledge, the first approach that integrates LLMs for program reduction task. LPR synergistically leverages the strengths of both language-generic program reducers and LLMs. Specifically, LPR alternates between invoking a language-generic reducer (we use Perses in experiments) and the LLM. Initially, Perses efficiently reduces large programs to a size manageable for the LLM. Subsequently, the LLM further transforms Perses’s output based on specific user-defined prompts that dictate the required transformations. Following this, Perses is re-invoked, as transformations made by the LLM often create additional opportunities for reduction. This process iterates until the program cannot be further minimized. For transformations, we have identified five language-generic transformations to enable further reduction: Function Inlining, Loop Unrolling, Data Type Elimination, Data Type Simplification, and Variable Elimination.
@@
We have conducted extensive evaluations on LPR, illustrating its superiority over Vulcan, the state-of-the-art language-generic algorithm. On three benchmark suites, i.e., Benchmark-C, Benchmark-Rust and Benchmark-JS, LPR produces significantly smaller programs than Vulcan by 24.93%, 4.47% and 11.71%, separately. Moreover, LPR and Vulcan complement each other to some extent. For the C language which C-Reduce is optimized for, by applying Vulcan to the output produced by LPR, we can attain program sizes that are on par with those achieved by C-Reduce. For efficiency, LPR performs comparably to Vulcan. In terms of execution time, LPR is more efficient than Vulcan on reducing complex programs. Furthermore, our detailed analysis indicates that each of the proposed transformations plays a crucial role in the reduction process. We also compare the performance with the multi-level prompt against that without it, illustrating the efficacy of our proposed multi-level prompting approach.
@@
We introduce LLMs-Aided Program Reduction (LPR), marking the first algorithm that integrates LLMs into the program reduction process. By synergizing the capabilities of both language-generic tools and LLMs, LPR achieves a balance between generality across various languages and awareness of semantics in specific languages. Simultaneously, it maintains flexibility in the design and extension of new transformations.
We propose a multi-level prompting approach to guide LLMs to execute program transformations, and demonstrate its effectiveness in practice. We propose five general-purposed transformations for LLMs to reduce programs or expose more reduction opportunities.
@@ 2.1.1. Language-Generic Reducers
Some reducers can be generalized to multiple languages. For instance, given the formal syntax of a programming language, algorithms like HDD and Perses can be used to reduce programs corresponding to that language. HDD parses the language into a parse tree and then applies the DDMin (Zeller and Hildebrandt, 2002) at each level of the tree to remove unnecessary tree nodes as much as possible. Perses goes further than HDD by performing certain transformations on the formal syntax to avoid generating syntactically incorrect program variants. Vulcan extends Perses, by introducing novel auxiliary reducers to exhaustively search for smaller variants by replacing identifiers/sub-trees and deleting local combinations of tree nodes on the parse tree.
However, different languages possess unique semantic features. Although the aforementioned algorithms are relatively efficient (Sun et al., 2018), they are incapable of utilizing unique semantics of a particular language to further reduce a program. For example, these algorithms lacks the ability to perform transformations like function inlining. Although Vulcan can identify more reduction opportunities through transformations such as identifier replacement and local exhaustive search, its approach is akin to ”brute force” enumeration. This method lacks awareness of the given program’s semantics, making it less effective and efficient overall.
@@ 2.1.2. Language-Specific Reducers
However, developing language-specific reducer is nontrivial. To the best of our knowledge, only a few languages have specific reducers, such as C (Regehr et al., 2012), Java (Kalhauge and Palsberg, 2019, 2021), and SMT-LIBv2 (Niemetz and Biere, 2013). The reason is that the process of designing a reducer for a language, or adding new transformations to an existing reducer, is a time-consuming and labor-intensive task. For instance, in version 2.10.0 of C-Reduce (Regehr, 2012), function inlining was implemented with 604 lines of C++ code. Such challenges impede the development and maintenance of language-specific reducers.
@@ 2.2. Large Language Models
Large Language Models refer to a type of deep learning technique that is trained on massively huge data sets for diverse tasks. The advent of LLMs has opened up numerous potential opportunities across diverse research fields. LLMs are not only proficient in processing natural languages but also exhibit substantial capabilities in understanding and processing programming languages. This highlights the promising future and evolutionary prospects in the realm of software engineering. Recently, LLMs have been applied and assessed on various software engineering tasks, such as automatic program repair (Huang et al., 2023; Xia and Zhang, 2023; Xia et al., 2023b, a; Fan et al., 2023) and program generation (Liu et al., 2023; Zhong and Wang, 2023; Tian et al., 2023a).
However, despite the usefulness of LLMs, some researchers (Li et al., 2023) illustrate that current LLMs are weak in distinguishing nuances between programs. Moreover, the memorizing and processing capacity of LLMs deteriorates as the input size grows, a.k.a., catastrophic forgetting (Kemker et al., 2018; Arora et al., 2019; Goodfellow et al., 2013). Moreover, one cannot expect LLMs to automatically complete complex tasks; they must be guided accordingly. Therefore, for program reduction, directly asking LLMs to reduce programs with tens of thousands of lines is impractical.
@@ 3. Approach
In this section, we first introduce a motivating example, then we provide an overview of the LPR workflow. We also outline the details of prompts and proposed transformations in the workflow that enable the LLM to function effectively with given programs.
@@ 3.1. Motivation
A motivating example is displayed in Figure 1, the original code contains highly nested loops, shown in Figure 0(a). From Figure 0(b) to Figure 0(c), the nested loops are fully unrolled into hundreds of lines via Loop Unrolling, based on the semantic transformations from the LLM. Despite the temporary size increase, the following Perses effectively eliminates all lines except for the bug-relevant one. This is also the final result of LPR, presented in Figure 0(d). By contrast, as shown in Figure 0(e), Vulcan is incapable of escaping the local minima by exhaustively replacing identifiers and tree nodes. Moreover, C-Reduce is not integrated with transformations to unroll loops, and thus cannot fully break down the loop structures. Even though loop unrolling techniques can be added into C-Reduce in future versions, it will be labor-intensive to implement a specific transformation compared to user-defined prompts in natural language.
@@ 3.2. Workflow
The overview of the workflow is outlined in Figure 2. Given a bug-triggering program as input, LPR invokes a language-generic reducer and the LLM alternately, until the target program cannot be further reduced. In each iteration, the language-generic reducer efficiently reduces the given program to 1-tree-minimality in syntax level. By contrast, the LLM leverages the semantic knowledge of the language and transforms the program given by the language-generic reducer. This process is guided by user-defined prompts, aiming to expose more reduction potentials to the language-generic reducer.
@@
Algorithm 1 shows LPR’s reduction algorithm. Given as inputs (1) a program targeted for reduction, (2) a property that must be preserved, and (3) the pre-defined prompts, LPR generates a reduced program . prompts contains primaryQuestion and followupQuestion. primaryQuestion instructs the LLM to identify a list of targets to be transformed, and followupQuestion guides the LLM to apply transformation on each target individually. They will be further introduced in section 3.4. Initially, LPR loads a sequence of transformations as delineated in Algorithm 1. It then iterates through each transformation, as detailed from Algorithm 1 to Algorithm 1. section 3.3 displays the details of each transformation.
@@
During this process, for each transformation, the algorithm retrieves a predefined primary question along with a follow-up question on Algorithm 1 – Algorithm 1. LPR first asks the LLM the primary question under the current program. This query aims to guide the LLM to generate a list of specific targets upon which the transformation will be executed. For instance, for Loop Unrolling in the motivating example Figure 0(b), the LLM is asked to identify a list of loops in the given program to be unrolled in the primaryQuestion, and returns a target list .
@@
On Algorithm 1 to 1, LPR uses followupQuestion to guide LLM to apply the transformation on each identified target within the program. In the motivating example, the followupQuestion can be framed as “Given the program { PROGRAM } and the loop for(i=0;...), optimize it via loop unrolling”. The modified program is then extracted from the LLM’s response text on Algorithm 1. In the example, all loops are unrolled into repeated lines of code in Figure 0(c).
@@
In experimental scenarios, given an input program and the prompt, the LLM may generate multiple transformed programs, as the number of responses can be customized. Among all transformed programs, LPR keeps the smallest one that still passes the property test, and discards others. If no transformed program returned from this query satisfies the property, LPR keeps the original one before this query. After each transformation is completed, language-agnostic reducers such as Perses (Sun et al., 2018) are employed to seek additional reduction opportunities, considering that the transformation might have introduced new potentials for further simplification. The algorithm persists in the outermost loop until it reaches a fixpoint, signifying that the program size can no longer be reduced.

### v3

@@ Abstract.
Our thorough evaluation on 50 benchmarks across three programming languages (i.e., C, Rust and JavaScript) has demonstrated LPR’s practicality and superiority over Vulcan, the state-of-the-art language-generic program reducer. For effectiveness, LPR surpasses Vulcan by producing 24.93%, 4.47%, and 11.71% smaller programs on benchmarks in C, Rust and JavaScript, separately. Moreover, LPR and Vulcan have the potential to complement each other. For the C language for which C-Reduce is optimized, by applying Vulcan to the output produced by LPR, we can attain program sizes that are on par with those achieved by C-Reduce. For efficiency perceived by users, LPR is more efficient when reducing large and complex programs, taking %, %, % less time than Vulcan to finish all the benchmarks in C, Rust and JavaScript, separately.
@@ 1. Introduction
Program reduction techniques (Sun et al., 2018; Xu et al., 2023; Regehr et al., 2012; Kalhauge and Palsberg, 2019, 2021; Niemetz and Biere, 2013; Tian et al., 2023c; Wang et al., 2021; Misherghi and Su, 2006; Donaldson et al., 2021) aim to facilitate compiler debugging by minimizing the bug-triggering programs with efficacy and efficiency. Given a program and a property that preserves, program reduction techniques (a.k.a., program reducers) produce a minimal program that still preserves . Program reduction has been widely used in various software engineering tasks (Donaldson and MacIver, 2021; Heo et al., 2018), especially in compiler testing and debugging (Livinskii et al., 2020; Wang et al., 2023a; Lecoeur et al., 2023; Tian et al., 2023b, d).
However, a critical challenge in program reduction has not been properly addressed, i.e., the trade-off between generality across languages and specificity to a certain language. Currently, there are two categories of program reduction techniques: language-specific reduction (Regehr et al., 2012; Kalhauge and Palsberg, 2019, 2021) and language-generic reduction (Zeller and Hildebrandt, 2002; Misherghi and Su, 2006; Sun et al., 2018; Xu et al., 2023; Zhang et al., 2023). The former category leverages language-specific semantics to transform and shrink programs in certain languages, while the latter only uses transformations applicable to any programming language. Although language-specific reducers are usually more effective in reduction, designing an effective reducer for a specific language, especially language-specific transformations, requires a deep understanding of language features and a significant amount of time and engineering effort. Therefore, only a limited set of languages have custom-designed reducers, such as C (Regehr et al., 2012), Java (Kalhauge and Palsberg, 2019, 2021), and SMT-LIBv2 (Niemetz and Biere, 2013). Meanwhile, language-generic reducers can be applied to diverse languages, but lack the knowledge of language features and semantics and thus are incapable of performing language-specific transformations (e.g., function inlining) that can enable further reduction. As a result, they cannot utilize peculiar features of each language to produce optimally reduced programs.
@@
Meanwhile, we also notice that recent Large Language Models (LLMs) could be powerful assistants in performing language-specific transformations, like how they perform in other software engineering tasks such as code generation and test generation (Sun et al., 2023; Jain et al., 2022; Wei et al., 2023; Deng et al., 2023; Fan et al., 2023a; Xia et al., 2023b; Gu, 2023). Specifically, LLMs are trained with massive programs, and they have started to exhibit the ability to analyze and transform programs in prevalent languages. If we can properly leverage this ability for program reduction, we may have a language-generic reducer being aware of the semantics of various prevalent languages. In addition, the utilization of LLMs can simplify the customization and extension of reducers, as it would be time-consuming and labor-intensive to manually implement a language-specific reducer or add functionality to an existing one (such as C-Reduce using the Clang frontend to implement C/C++-specific program transformations).
Challenges of Using LLMs. LLMs are not the silver bullet to program reduction. When LLMs are instructed to handle source code as inputs, due to the intrinsic limitations of LLMs (Fan et al., 2023b; Wang et al., 2024; Li et al., 2023a), LLMs may be unable to understand subtle differences in code (Li et al., 2023b) and be distracted by irrelevant context (Shi et al., 2023). Moreover, in program reduction, input programs typically contain tens of thousands of line of code, which surpasses the input limits of LLMs. Besides, without effective guidance, LLMs are unclear about what transformations to perform.
LLMs-Aided Program Reduction (LPR). We propose LPR in this paper, which is, to the best of our knowledge, the first approach that integrates LLMs for program reduction tasks. LPR synergistically leverages the strengths of both language-generic program reducers and LLMs. Specifically, LPR alternates between invoking a language-generic reducer (we use Perses in experiments) and the LLM. Initially, Perses efficiently reduces large programs to a size manageable for the LLM. Subsequently, the LLM further transforms Perses’s output based on specific user-defined prompts that dictate the required transformations. Following this, Perses is re-invoked, as transformations made by the LLM often create additional opportunities for reduction. This process iterates until the program cannot be further minimized. For transformations, we have identified five language-generic transformations to enable further reduction: Function Inlining, Loop Unrolling, Data Type Elimination, Data Type Simplification, and Variable Elimination.
@@
We have conducted extensive evaluations on LPR, illustrating its superiority over Vulcan, the state-of-the-art language-generic algorithm. On three benchmark suites, i.e., Benchmark-C, Benchmark-Rust and Benchmark-JS, LPR produces significantly smaller programs than Vulcan by 24.93%, 4.47% and 11.71%, separately. Moreover, LPR and Vulcan complement each other. For C language which C-Reduce is optimized for, by applying Vulcan to the output produced by LPR, we attain program sizes that are on par with those achieved by C-Reduce. For reduction efficiency, LPR performs comparably to Vulcan. In terms of execution time perceived by users, LPR is more efficient than Vulcan on reducing complex programs. Furthermore, our detailed analysis indicates that each of the proposed transformations plays a crucial role in the reduction process. We also compare the LPR’s performance with the multi-level prompt against that without it, illustrating the efficacy of our proposed multi-level prompting approach.
@@
We introduce LPR, the first attempt to use LLMs for program reduction. By synergizing the capabilities of both language-generic program reducers and LLMs, LPR exhibits both generality across various languages and awareness of semantics in specific languages. Moreover, LPR is easy and flexible to extend with new transformations by simply designing new prompts.
We propose a multi-level prompting approach to guide LLMs to execute program transformations, and demonstrate its effectiveness in practice. We propose five general-purposed transformations for LLMs to reduce programs or create more reduction opportunities for language-generic program reducers to exploit.
@@ 2.1.1. Language-Generic Reducers
Some reducers can be generalized to multiple languages. For instance, given the formal syntax of a programming language, algorithms like HDD and Perses can be used to reduce programs corresponding to this language. HDD parses the program into a parse tree and then applies the DDMin (Zeller and Hildebrandt, 2002) at each level of the tree to remove unnecessary tree nodes as much as possible. Perses goes further than HDD by performing certain transformations on the formal syntax to avoid generating syntactically incorrect program variants. Vulcan extends Perses, by introducing novel auxiliary reducers to exhaustively search for smaller variants by replacing identifiers/sub-trees and deleting local combinations of tree nodes on the parse tree.
However, different languages possess their own unique semantic features. Although the aforementioned algorithms are relatively efficient (Sun et al., 2018), they are incapable of utilizing unique semantics of a particular language to further reduce a program. For example, these algorithms lacks the ability to perform transformations like function inlining. Although Vulcan can identify more reduction opportunities through transformations such as identifier replacement and local exhaustive search, its approach is akin to ”brute force” enumeration. This method lacks awareness of the given program’s semantics, making it less effective and efficient overall.
@@ 2.1.2. Language-Specific Reducers
However, developing a language-specific reducer is nontrivial. To the best of our knowledge, only a few languages have specific reducers, such as C (Regehr et al., 2012), Java (Kalhauge and Palsberg, 2019, 2021), and SMT-LIBv2 (Niemetz and Biere, 2013). The reason is that the process of designing a reducer for a language, or adding new transformations to an existing reducer, is time-consuming and labor-intensive. For instance, in version 2.10.0 of C-Reduce (Regehr, 2012), function inlining was implemented with 604 lines of C++ code. Such challenges impede the development and maintenance of language-specific reducers.
@@ 2.2. Large Language Models
Large Language Models refer to a type of deep learning models trained on huge data sets for diverse tasks. The advent of LLMs has opened up numerous potential opportunities across diverse research fields. LLMs are not only proficient in processing natural languages but also exhibit substantial capabilities in understanding and processing programming languages. This highlights the promising future and evolutionary prospects in the realm of software engineering. Recently, LLMs have been applied and assessed on various software engineering tasks, such as automatic program repair (Huang et al., 2023; Xia and Zhang, 2023; Xia et al., 2023b, a; Fan et al., 2023a) and program generation (Liu et al., 2023a; Zhong and Wang, 2023; Tian et al., 2023a).
However, despite the usefulness of LLMs, some researchers (Li et al., 2023b) illustrate that current LLMs are weak in distinguishing nuances between programs. Moreover, the memorizing and processing capacity of LLMs deteriorates as the input size grows (Fan et al., 2023b; Shaham et al., 2022). Besides, one cannot expect LLMs to automatically complete complex tasks; they must be guided accordingly (Xia et al., 2023b; Li et al., 2023b). Therefore, for program reduction, directly asking LLMs to reduce programs with tens of thousands of lines is impractical.
@@ 3. Approach
In this section, we first introduce a motivating example, and then we provide an overview of the LPR workflow. We also outline the details of prompts and proposed transformations in the workflow that enable the LLM to function effectively on given programs.
@@ 3.1. Motivation
A motivating example is displayed in Figure 1, where the original code contains highly nested loops, shown in Figure 1(a). From Figure 1(b) to Figure 1(c), the nested loops are fully unrolled into hundreds of lines via Loop Unrolling, based on the semantic transformations from the LLM. Despite the temporary size increase, the following Perses effectively eliminates all lines except for the bug-relevant one. This is also the final result of LPR, presented in Figure 1(d). By contrast, as shown in Figure 1(e), Vulcan is incapable of escaping the local minima by exhaustively replacing identifiers and tree nodes. Similarly, C-Reduce cannot fully break down the loop structures, given that it is not integrated with transformations to unroll loops. Even though loop unrolling techniques can be added into C-Reduce in future versions, it will be labor-intensive to implement a specific transformation compared to user-defined prompts in natural language.
@@ 3.2. Workflow
The overview of the workflow is outlined in Figure 2. Given a bug-triggering program as input, LPR invokes a language-generic reducer and the LLM alternately, until the target program cannot be further reduced. In each iteration, the language-generic reducer efficiently reduces the given program to a minimal one. By contrast, the LLM leverages the semantic knowledge of the language and transforms the program given by the language-generic reducer. This process is guided by user-defined prompts, aiming to expose more reduction potentials to the language-generic reducer.
@@
Algorithm 1 shows LPR’s reduction algorithm. Given as inputs (1) a program targeted for reduction, (2) a property that must be preserved, and (3) the pre-defined prompts, LPR generates a reduced program . The pre-defined prompts consist of primaryQuestion and followupQuestion, in which primaryQuestion instructs the LLM to identify a list of targets to be transformed, and followupQuestion guides the LLM to apply transformation on each target individually. They will be further introduced in section 3.4. Initially, LPR calls Perses to quickly minimize , so that the size of the program becomes manageable for the LLM to process. Next, LPR loads a sequence of transformations as delineated in Algorithm 1, and then iterates through each transformation, as detailed from Algorithm 1 to Algorithm 1. section 3.3 displays the details of each transformation.
@@
During this process, for each transformation, the algorithm retrieves a predefined primary question along with a follow-up question on Algorithm 1 – Algorithm 1. LPR firstly asks the LLM the primary question under the current program. This query aims to guide the LLM to generate a list of specific targets upon which the transformation will be executed. For instance, for Loop Unrolling in the motivating example Figure 1(b), the LLM is asked to identify a list of loops in the given program to be unrolled in the primaryQuestion, and returns a target list .
@@
On Algorithm 1 to 1, LPR uses followupQuestion to guide the LLM to apply the transformation on each identified target within the program. In the motivating example, the followupQuestion can be framed as “Given the program { PROGRAM } and the loop for(i=0;...), optimize it via loop unrolling”. The modified program is then extracted from the LLM’s response text on Algorithm 1. In the example, all loops are unrolled into repeated lines of code in Figure 1(c).
@@
In experimental scenarios, given an input program and the prompt, the LLM may generate multiple transformed programs, as the number of responses can be customized. Among all transformed programs, LPR keeps the smallest one that still passes the property test, and discards others. If no transformed program returned from this query satisfies the property, LPR keeps the original one before this query. After each transformation is completed, a language-generic reducer such as Perses (Sun et al., 2018) is employed to seek additional reduction opportunities, considering that the transformation might have introduced new potential for further simplification. The algorithm persists in the outermost loop until it reaches a fixpoint, signifying that the program size can no longer be reduced.